OpenStudy (anonymous):

MEDEAL AND FAN Find the sum or difference. 1. Mr.Jensen bought 2 2/3 gallons of gasoline yesterday. Today he bought 3 1/2 more gallons. How many total gallons of gasoline has he bought? A 5 3/5 gallons B 5 3/6 gallons C 6 1/6 gallons D 5 1/6 gallons

jimthompson5910 (jim_thompson5910):

You need to add up \(\Large 2 \frac{2}{3}\) and \(\Large 3 \frac{1}{2}\). Don't forget to get the lowest common denominator (LCD).

jimthompson5910 (jim_thompson5910):

In this case, the LCD is 6 because 3*2 = 6. I'm multiplying the denominators.

OpenStudy (anonymous):

so c